# Resisting Contextual Interference in RAG via Parametric-Knowledge Reinforcement

### v3

@@ Abstract
Retrieval-augmented generation (RAG) improves performance on knowledge-intensive tasks but can be derailed by wrong, irrelevant, or conflicting retrieved text, causing models to rely on inaccurate evidence and cascade errors. We propose Knowledgeable-R1, a reinforcement-learning framework that explicitly trains large language models to use parametric knowledge (PK) to resist contextual interference while still exploiting external context when it is reliably helpful. Knowledgeable-R1 introduces a joint sampling scheme that generates paired responses with and without retrieval, and learns both local advantages (within each decoding regime) and global advantages under the same input to quantify when to ignore misleading context versus adopt it. We employ an asymmetric advantage transformation that amplifies exploratory behaviors toward parametric knowledge. Experiments show that Knowledgeable-R1 significantly improves robustness and reasoning accuracy in knowledge conflict scenarios and general RAG scenarios, outperforming SOTA baselines by 23% in counterfactual scenarios, and without degradation when the retrieved context is fully accurate. Our code are available at https://github.com/lcy80366872/knowledgeable-R1.
@@ 1 Introduction
Figure 1: Our approach combines parametric (blue) and context-based (pink) path sampling with policy optimization to ensure correct outputs, even when erroneous context inputs. Green blocks represent positive answers, while red blocks represent negative ones.
The goal of this work is to develop a reinforcement learning framework that explores both contextual and parametric knowledge after retrieval input. We call our method Knowledgeable-R1 because it is a knowledge-aware reasoning approach that uses parametric knowledge to mitigate contextual interference in the RAG system. A conceptual overview of our framework is presented in Figure 1. Specifically, Knowledgeable-R1 enables LLMs to (1) explore both parametric and contextual knowledge through joint sampling, (2) distinguish the relative merits of these two knowledge types using locally and globally defined advantages, and (3) mitigate penalties for less likely parametric outputs by applying an adaptive asymmetric advantages transformation. We conducted experiments across five scenarios, demonstrating that Knowledgeable-R1 outperforms a range of RAG approaches by a significant margin. By effectively utilizing both parametric and contextual knowledge, Knowledgeable-R1 shows significant improvements in adversarial context settings, achieving a 23% improvement over GRPO baselines and a 30.47% improvement over original RAG prompting, while maintaining strong performance in reliable contexts.
@@ 3.3 Advantage Calculation with Local and Global Normalization
$A^{\mathrm{pk\mbox{-}local}}_{i}=\frac{R^{\mathrm{pk}}_{i}-\mathrm{mean}(\{R^{\mathrm{pk}}_{k}\})}{\mathrm{std}(\{R^{\mathrm{pk}}_{k}\})+\varepsilon}.$
@@
$A^{\mathrm{ck\mbox{-}local}}_{j}=\frac{R^{\mathrm{ck}}_{j}-\mathrm{mean}(\{R^{\mathrm{ck}}_{k}\})}{\mathrm{std}(\{R^{\mathrm{ck}}_{k}\})+\varepsilon},\quad A^{\mathrm{ck\mbox{-}global}}_{j}=\frac{R^{\mathrm{ck}}_{j}-\mathrm{mean}(\mathcal{U}_{p^{\prime}})}{\mathrm{std}(\mathcal{U}_{p^{\prime}})+\varepsilon}.$
@@ 3.4 Knowledge Balance Modulation
$\beta\leftarrow\mathrm{clip}\left(\frac{S_{\mathrm{ck}}-S_{\mathrm{rpk}+}}{S_{\mathrm{rpk}-}},0.01,1\right),$
@@ 3.5 Policy Optimization
$J_{\mathrm{PK}}=\frac{1}{n_{\text{pk}}}\sum_{i=1}^{n_{\text{pk}}}\sum_{t=1}^{|o_{i}|}\min\Big[r^{\mathrm{PK}}_{i,t}\,A_{i},\;\mathrm{clip}(r^{\mathrm{PK}}_{i,t},\,1-\epsilon,\,1+\epsilon)\,A_{i}\Big],$
@@
$J_{\mathrm{RPK}}=\frac{1}{n_{\text{pk}}}\sum_{i=1}^{n_{\text{pk}}}\sum_{t=1}^{|o_{i}|}\min\Big[r^{\mathrm{RPK}}_{i,t}\,T(\hat{A}_{i}),\;\mathrm{clip}(r^{\mathrm{RPK}}_{i,t},\,1-\epsilon,\,1+\epsilon)\,T(\hat{A}_{i})\Big],$
@@
The goal of our method is to optimize $\theta$ to maximizes $\mathcal{J}(\theta)$, the full training procedure is outlined in Algorithm 1 in Appendix E.
@@ 4.1.1 Reasoning Instruction and Baselines
We evaluate with Qwen2.5-3B-Instruct, Qwen2.5-7B-Instruct, Qwen2.5-14B-Instruct and Llama3.1-8B-Instruct. Baselines include query-only prompting, RAG prompting, Astute-RAG (a prompting method for imperfect retrieval (Wang et al., 2025a)), CK-PLUG(Bi et al., 2025c) and GRPO(Guo et al., 2025a) with RAG. We evaluate using exact match (EM) (Jin et al., 2025). To ensure fairness, all methods use the same total training data and are trained for one epoch. For GRPO and Knowledgeable-R1, we use a global batch size of 128, a rollout batch size of 32, a rollout temperature of 1, and a learning rate of $1\times 10^{-6}$. All experiments run on 8 H100 GPUs, using the same system prompt for both training and inference, where the inference input combines the query and context. For further experiment details, see Appendix B, C, and D.
@@ 4.1.2 Datasets
We introduce KnowQA to test methods across the five scenarios. First, we test parametric knowledge using query-only prompts (Appendix A), where a correct answer means the model has the relevant PK. Then, we create contextual inputs using the top-5 retrieved paragraphs. This method is suitable for knowledge-intensive tasks. KnowQA includes scientific, factual, and commonsense questions, organized by scenario:
@@
  • Irrelevant Context. We use ExplainCPE (Wen et al., 2024), a medical QA dataset from the Chinese National Licensed Pharmacist Examination. It tests the impact of knowledge mismatches on model performance, where adding retrieved documents can lower accuracy compared to query-only.
@@ 4.2 Experimental Results
We evaluate Knowledgeable-R1 in various contexts and find that it effectively selects reliable evidence while ignoring misleading or irrelevant information, especially when parametric knowledge (PK) is sufficient. This leads to significant improvements in challenging interference settings (S2-S5) without harming performance in correct contexts (S1). Consistent results are observed across models of different sizes (3B, 7B, 8B, 14B), with detailed results for Qwen2.5-7B-Instruct and Llama3.1-8B-Instruct in Tables 2 and 3, and similar patterns for other models (see Appendix F). Our experimental analysis highlights three key strengths of Knowledgeable-R1:
@@ 4.2.2 Analysis on Parametric-Knowledge Answerable Subset
We define the parametric-knowledge answerable subset as questions where the model provides correct answers using query-only prompting, showing that it has enough parametric knowledge. When focusing on this subset, all methods improve, but Knowledgeable-R1 shows the biggest gains when the context is wrong or mixed (Table 3).
@@
Method PC-MR PC-MC PC-QA NC-MR NC-MC NC-QA SC ExplainPE HotPotQA
2Wiki
MultiHopQA
Musique
@@
Method PC-MR PC-MC PC-QA NC-MR NC-MC NC-QA SC ExplainPE HotPotQA
2Wiki
MultiHopQA
Musique
@@
On Qwen2.5-7B-Instruct, Knowledgeable-R1 achieves 89.61%/79.62%/66.00% on NC-MR/MC/QA, greatly outperforming RAG prompting (32.47%/19.75%/28.50%) and GRPO w/ RAG (57.61%/46.50%/62.50%). Similar results are observed for Llama3.1-8B-Instruct. These results show that Knowledgeable-R1 effectively learns when parametric knowledge is reliable and ignores conflicting context. In correct contexts (S1), Knowledgeable-R1 maintains competitive performance.
The most impressive gains appear in S5 (partially relevant context), where Knowledgeable-R1 achieves 78.49% compared to 66.54% for GRPO w/ RAG on HotpotQA with Qwen2.5-7B-Instruct, and 83.55% compared to 73.17% on Llama3.1-8B-Instruct. This shows that the learned policy effectively identifies useful evidence, keeping relevant information and ignoring distractions.
@@ 4.2.3 Ablation Studies
We conduct ablation studies on the Qwen2.5-7B-Instruct model to evaluate the contribution of each Knowledgeable-R1 component (Table 4). The largest performance drops occur in TIFE (parametric knowledge correct, context knowledge incorrect) scenarios, highlighting the importance of each component in handling misleading context.
@@
We assess the impact of three policy types: PK (parametric knowledge), CK (context-aware), and RPK (robust parametric knowledge). Removing the parametric knowledge objective ($J_{PK}$) results in moderate performance drops, particularly in FIFE scenarios (-1.11% for MC, -2.14% for QA). The most significant degradation occurs when the relative parametric knowledge reward ($J_{RPK}$) is removed, causing catastrophic performance losses in TIFE scenarios (-33.12% for MC, -3.50% for QA). This demonstrates the crucial role of the RPK policy in ensuring reliable reasoning under contextual interference.
@@
Removing the local context advantage ($A^{ck-local}$) causes a significant performance drop in context-answerable scenarios, while removing the global context advantage ($A^{ck-global}$) reduces TIFE performance. This suggests that local advantages boost the performance of their respective types, such as CK samples performing better when the context is correct. Local advantages also extend the influence of their type over others. These findings highlight the importance of local and global advantages in optimizing CK and RPK policies.
@@
The adaptive $\beta$ modulation is vital for maintaining parametric knowledge as a fallback option. Using fixed $\beta$ values reduces TIFE performance by 27.39% for MC and 12.50% for QA, demonstrating the importance of dynamic balancing between knowledge sources. The substantial performance drop in FIFE scenarios further highlights the effectiveness of adaptive balancing, particularly when both knowledge sources are unreliable. This supports our design of reducing penalties for RPK when parametric knowledge is slightly less effective than context, ensuring parametric knowledge remains usable in noisy contexts.
These results affirm that Knowledgeable-R1’s components work synergistically to balance contextual and parametric knowledge, enabling robust reasoning in challenging scenarios.
@@ 5 Conclusion
We introduce Knowledgeable-R1, a reinforcement learning framework that uses joint sampling to balance contextual and parametric knowledge in large language models (LLMs). Our evaluation across five scenarios—correct, adversarial, self-conflicting, irrelevant, and partially relevant—shows significant improvements, especially in adversarial settings, while maintaining strong performance in reliable contexts. Future work will focus on expanding Knowledgeable-R1 to more complex environments, like multi-model RAG systems and integrating multiple exploration cues into a mixed objective-conditioned policy.
@@
To facilitate the reproducibility of our work, we have made the following available: (1) the source code of our algorithm, which conducts training using the VERL framework (Sheng et al., 2025), along with supplementary materials; (2) complete experimental configurations in C. Additionally, a detailed account of the data pre-processing procedures for the datasets in Section 4.1.2 can be found in C. We also plan to release the full open-source code in the future.
@@
Appendix A Prompt designed in Knowledgeable-R1 and baselines

### v4

@@ Abstract
Retrieval-augmented generation (RAG) improves performance on knowledge-intensive tasks but can be derailed by wrong, irrelevant, or conflicting retrieved text, causing models to rely on inaccurate evidence and cascade errors. We propose Knowledgeable-R1, a reinforcement-learning framework that explicitly trains large language models to use parametric knowledge (PK) to resist contextual interference while still exploiting external context when it is reliably helpful. Knowledgeable-R1 introduces a joint sampling scheme that generates paired responses with and without retrieval, and learns both local advantages (within each decoding regime) and global advantages under the same input to quantify when to ignore misleading context versus adopt it. We employ an asymmetric advantage transformation that amplifies exploratory behaviors toward parametric knowledge. Experiments show that Knowledgeable-R1 significantly improves robustness and reasoning accuracy in knowledge conflict scenarios and general RAG scenarios, outperforming SOTA baselines by +22.89% in counterfactual scenarios, and without degradation when the retrieved context is fully accurate. Our code are available at https://github.com/lcy80366872/knowledgeable-R1.
@@ 1 Introduction
Figure 1: Overview of Knowledgeable-R1, which integrates parametric (blue) and contextual (pink) sampling with policy optimization to ensure robust reasoning under misleading context. Green and red blocks indicate correct and incorrect responses, respectively. The inset chart displays accuracy on the ConFiQA-MC adversarial scenario (S2), where Knowledgeable-R1 (37.3%) significantly outperforms GRPO w/ RAG (19.7%), query-only prompting (25.8%), and RAG prompting (8.1%); see Table 2.
The goal of this work is to develop a reinforcement learning framework that explores both contextual and parametric knowledge after retrieval input. We call our method Knowledgeable-R1 because it is a knowledge-aware reasoning approach that uses parametric knowledge to mitigate contextual interference in the RAG system. A conceptual overview of our framework is presented in Figure 1. Specifically, Knowledgeable-R1 enables LLMs to (1) explore both parametric and contextual knowledge through joint sampling, (2) distinguish the relative merits of these two knowledge types using locally and globally defined advantages, and (3) mitigate penalties for less likely parametric outputs by applying an adaptive asymmetric advantages transformation. We conducted experiments across five scenarios, demonstrating that Knowledgeable-R1 outperforms a range of RAG approaches by a significant margin. By effectively utilizing both parametric and contextual knowledge, Knowledgeable-R1 shows significant improvements in adversarial context settings, achieving +22.9% improvement over GRPO baselines in counterfactual scenarios (Table 3) and a 30.47% improvement over original RAG prompting, while maintaining strong performance in reliable contexts.
@@ 3.1 Task Definition
Clarification on RPK sampling. RPK does not generate an independent answer. For each query, we first sample a PK trajectory $o^{\mathrm{pk}}=(o^{\mathrm{pk}}_{1},\dots,o^{\mathrm{pk}}_{T})$ from $\pi_{\theta}(\cdot\mid p,\cdot)$ (query-only). In the RPK branch, the same token sequence $o^{\mathrm{pk}}$ is used as the target, but evaluated under the query+context prompt $p^{\prime}$: at each step $t$, we compute $\pi_{\theta}(o^{\mathrm{pk}}_{t}\mid p^{\prime},o^{\mathrm{pk}}_{<t})$ and maximize its log-probability under the RPK reward. In other words, RPK re-evaluates the PK answer under $p^{\prime}$ and encourages the model to maintain PK tokens even when misleading context is present.
@@ 3.3 Advantage Calculation with Local and Global Normalization
$A^{\mathrm{pk\mbox{-}local}}_{i}=\frac{R^{\mathrm{pk}}_{i}-\mathrm{mean}(\{R^{\mathrm{pk}}_{i}\})}{\mathrm{std}(\{R^{\mathrm{pk}}_{i}\})+\varepsilon}.$
@@
$A^{\mathrm{ck\mbox{-}local}}_{j}=\frac{R^{\mathrm{ck}}_{j}-\mathrm{mean}(\{R^{\mathrm{ck}}_{j}\})}{\mathrm{std}(\{R^{\mathrm{ck}}_{j}\})+\varepsilon},\quad A^{\mathrm{ck\mbox{-}global}}_{j}=\frac{R^{\mathrm{ck}}_{j}-\mathrm{mean}(\mathcal{U}_{p^{\prime}})}{\mathrm{std}(\mathcal{U}_{p^{\prime}})+\varepsilon}.$
@@ 3.4 Knowledge Balance Modulation
$\vskip-5.0pt\beta\leftarrow\mathrm{clip}\left(\frac{S_{\mathrm{ck}}-S_{\mathrm{rpk}+}}{S_{\mathrm{rpk}-}},0.01,1\right),\vskip-5.0pt$
@@ 3.5 Policy Optimization
$J_{\mathrm{PK}}$ $=\frac{1}{n_{\text{pk}}}\sum_{i=1}^{n_{\text{pk}}}\sum_{t=1}^{|o_{i}|}\min\Big[r^{\mathrm{PK}}_{i,t}\,A_{i},\;\mathrm{clip}(r^{\mathrm{PK}}_{i,t},1-\epsilon,1+\epsilon)\,A_{i}\Big],$ (1)
@@
$J_{\mathrm{RPK}}$ $=\frac{1}{n_{\text{pk}}}\sum_{i=1}^{n_{\text{pk}}}\sum_{t=1}^{|o_{i}|}\min\Big[r^{\mathrm{RPK}}_{i,t}\,T(\hat{A}_{i}),\;\mathrm{clip}(r^{\mathrm{RPK}}_{i,t},1-\epsilon,1+\epsilon)\,T(\hat{A}_{i})\Big].$
The total objective is the weighted sum of the individual objectives:
$\mathcal{J}(\theta)=\lambda_{\mathrm{pk}}J_{\mathrm{PK}}+\lambda_{\mathrm{ck}}J_{\mathrm{CK}}+\lambda_{\mathrm{rpk}}J_{\mathrm{RPK}}$ (2)
@@
The goal of our method is to optimize $\theta$ to maximizes $\mathcal{J}(\theta)$, the full training procedure is outlined in Algorithm 1 in Appendix L.
@@ 4.1.1 Reasoning Instruction and Baselines
We evaluate with Qwen2.5-3B-Instruct, Qwen2.5-7B-Instruct, Qwen2.5-14B-Instruct and Llama3.1-8B-Instruct. Baselines include query-only prompting, RAG prompting, Astute-RAG (a prompting method for imperfect retrieval (Wang et al., 2025a)), CK-PLUG (Bi et al., 2025c), SFT, and GRPO (Guo et al., 2025a) with RAG. We evaluate using exact match (EM) (Jin et al., 2025). To ensure fairness, all methods use the same total training data and are trained for one epoch. For GRPO and Knowledgeable-R1, we use a global batch size of 128, a rollout batch size of 32, a rollout temperature of 1, and a learning rate of $1\times 10^{-6}$. All experiments run on 8 H100 GPUs, using the same system prompt for both training and inference, where the inference input combines the query and context. For further experiment details, see Appendix I, J, and K.
Baseline clarifications. CK-PLUG (Bi et al., 2025c) is a plug-and-play decoding method that detects knowledge conflicts via confidence gain and adjusts token probabilities for conflict spans, enabling fine-grained control over parametric vs. contextual reliance. GRPO w/ RAG trains only the CK objective with query+context rollouts (no PK/RPK branches), using the same number of total rollouts as Knowledgeable-R1. SFT is a strong fine-tuning baseline trained on the same mixture of good/bad context data as Knowledgeable-R1, with explicit conflict-handling instructions. Our method splits rollouts into PK, CK, and RPK groups; roughly half the rollout budget is allocated to query-only PK trajectories, making the two methods directly comparable.
@@ 4.1.2 Datasets
We introduce KnowQA to test methods across the five scenarios. First, we test parametric knowledge using query-only prompts (Appendix H), where a correct answer means the model has the relevant PK. Then, we create contextual inputs using the top-5 retrieved paragraphs. This method is suitable for knowledge-intensive tasks. KnowQA includes scientific, factual, and commonsense questions, organized by scenario:
@@
  • Irrelevant Context. We use ExplainPE (Wen et al., 2024), a medical QA dataset from the Chinese National Licensed Pharmacist Examination. It tests the impact of knowledge mismatches on model performance, where adding retrieved documents can lower accuracy compared to query-only.
@@
4.1.3 Retrieval Setup
We use dataset-provided contexts without any additional retriever. The contexts (including both correct and incorrect passages) are directly taken from the benchmark construction, without re-indexing or re-retrieval from a larger corpus. We adopt the top-5 paragraphs provided by these benchmarks, preserve their original order, and only truncate when the backbone model’s maximum length is exceeded. For ConFiQA training, correct and incorrect contexts are mixed in equal proportion.
@@ 4.2 Experimental Results
We evaluate Knowledgeable-R1 in various contexts and find that it effectively selects reliable evidence while ignoring misleading or irrelevant information, especially when parametric knowledge (PK) is sufficient. This leads to significant improvements in challenging interference settings (S2-S5) without harming performance in correct contexts (S1). Consistent results are observed across models of different sizes (3B, 7B, 8B, 14B), with detailed results for Qwen2.5-7B-Instruct and Llama3.1-8B-Instruct in Tables 2 and 3, and similar patterns for other models (see Appendix M). Our experimental analysis highlights three key strengths of Knowledgeable-R1:
@@ 4.2.2 Analysis on Parametric-Knowledge Answerable Subset
We define the parametric-knowledge answerable subset as questions where the model provides correct answers using query-only prompting, showing that it has enough parametric knowledge. When focusing on this subset, all methods improve, but Knowledgeable-R1 shows the biggest gains when the context is wrong or mixed (Table 3). Specifically, on NC-MR, NC-MC, and NC-QA, our method achieves an average improvement of 22.89% over GRPO w/ RAG.
@@
Method PC-MR PC-MC PC-QA NC-MR NC-MC NC-QA SC ExplainPE HotPotQA 2Wiki MultiHopQA Musique
@@
Method PC-MR PC-MC PC-QA NC-MR NC-MC NC-QA SC ExplainPE HotPotQA 2Wiki MultiHopQA Musique
@@
For Qwen2.5-7B-Instruct, Knowledgeable-R1 achieves 89.61%/79.62%/66.00% on NC-MR/MC/QA, significantly outperforming RAG prompting (32.47%/19.75%/28.50%) and GRPO w/ RAG (57.61%/46.50%/62.50%). Consistent trends are observed for Llama3.1-8B-Instruct. These results confirm that our method effectively identifies when parametric knowledge is reliable, allowing it to ignore conflicting information while maintaining stable performance in correct contexts (S1).
Significant improvements are observed in Scenario V (partially relevant context). On HotpotQA, Knowledgeable-R1 achieves 78.49% vs. 66.54% (GRPO w/ RAG) using Qwen2.5-7B-Instruct, and 83.55% vs. 73.17% using Llama3.1-8B-Instruct. This indicates that the learned policy effectively identifies useful evidence while filtering out distractions.
@@ 4.2.3 Ablation Studies
We conduct ablation studies on Qwen2.5-7B-Instruct to evaluate the contribution of each component (Table 4). The largest performance drops occur in TIFE (parametric correct, context incorrect) scenarios, highlighting the importance of each component in handling misleading context.
@@
We assess PK, CK, and RPK policies. Removing the parametric objective ($J_{PK}$) results in moderate drops, particularly in FIFE scenarios (-1.11% for MC, -2.14% for QA). The most significant degradation occurs when the relative parametric reward ($J_{RPK}$) is removed, causing catastrophic losses in TIFE scenarios (-33.12% for MC, -3.50% for QA). This confirms $J_{RPK}$ is crucial for robustness under contextual interference.
@@
Removing $A^{ck-local}$ causes significant performance drops in context-answerable scenarios, while removing $A^{ck-global}$ reduces TIFE performance. This suggests local advantages boost respective types (e.g., CK performs better when context is correct), while global advantages optimize the overall knowledge balance. Further ablation in Appendix B shows the framework is robust to the local-global ratio choice.
@@
Adaptive $\beta$ modulation is vital; fixed $\beta$ reduces TIFE performance by 27.39% (MC) and 12.50% (QA). The drop in FIFE scenarios further highlights the effectiveness of adaptive balancing when both knowledge sources are unreliable. Together, these components work synergistically to enable robust reasoning in challenging scenarios.
4.3 Additional Analysis
In this section, we investigate performance trade-offs, robustness, and the dynamic behavior of our framework.
Knowledge Balance. A critical component of Knowledgeable-R1 is the adaptive $\beta$ modulation. As shown in Table 5, the adaptive balancing scheme significantly outperforms any single fixed $\beta$ value across different datasets without requiring per-dataset tuning. Fixed values often create a trade-off: a low $\beta$ might favor parametric knowledge but hurt contextual reasoning, while a high $\beta$ does the opposite. Dynamically adjusting this balance based on prediction confidence ensures robust handling of conflicting information. Furthermore, during training, $\beta$ rapidly converges to a stable regime (stabilizing at 0.01 within 8 steps on ConFiQA, see Appendix A), indicating that the policy quickly learns to shield parametric knowledge from unreliable contexts.
Table 5: $\beta$ sensitivity study: fixed vs. adaptive. Adaptive $\beta$ achieves best-or-near-best across tasks without per-dataset tuning.
$\beta$ MC-avg QA-avg
0.01 56.03% 52.83%
0.2 52.74% 52.53%
0.5 51.19% 52.63%
0.8 48.22% 51.75%
1.0 48.05% 52.52%
Adaptive $\beta$ (ours) 56.43% 54.85%
Robustness to Data and Weights. A common concern is whether the model overfits to the "noise ratio" of the training data. However, even when trained on 99% correct context (only 1% bad), Knowledgeable-R1 still outperforms GRPO on S2 scenarios (Appendix E), proving it learns a genuine decision boundary rather than relying on dataset statistics. Additionally, regarding the S1 (correct context) performance gap compared to GRPO w/ RAG, ablations show this is tunable via the $J_{\mathrm{CK}}$ weight. Increasing its weight (e.g., ratio 1:2:1) closes the S1 gap while preserving most S2 advantages (Appendix C), demonstrating a controllable Pareto front between contextual compliance and parametric safety.
@@ 5 Conclusion
In this work, we presented Knowledgeable-R1, a principled reinforcement learning framework that systematically addresses the challenges of knowledge integration in retrieval-augmented generation. By employing a joint sampling strategy and asymmetric policy optimization, our approach enables models to effectively distinguish between reliable external context and internal parametric knowledge. Extensive benchmarking across diverse scenarios highlights that Knowledgeable-R1 achieves superior robustness against contextual interference, particularly in adversarial and conflicting settings, while preserving competitive performance on standard RAG tasks. These results underscore the potential of reinforcement learning for calibrating knowledge utilization. Future work will explore the scalability of this approach to more complex multi-source retrieval environments and investigate the integration of diverse exploration cues to further enhance model reasoning and grounding.
@@
To facilitate the reproducibility of our work, we have made the following available: (1) the source code of our algorithm, which conducts training using the VERL framework (Sheng et al., 2025), along with supplementary materials; (2) complete experimental configurations in J. Additionally, a detailed account of the data pre-processing procedures for the datasets in Section 4.1.2 can be found in J. We also plan to release the full open-source code in the future.
@@
Appendix A Training Dynamics of $\beta$
In our adaptive scheme, $\beta$ acts as a dynamic gate that modulates the influence of the relative parametric reward. As shown in Table 6, on the ConFiQA-MC dataset, $\beta$ undergoes a rapid descent during the initial stages of training, dropping from $0.23$ and stabilizing at a very low value ($0.01$). This rapid convergence is significant: it indicates that the reinforcement learning process quickly identifies the inherent unreliability of the external context in adversarial scenarios. By lowering $\beta$, the model effectively increases the penalty for deviating from its correct internal parametric knowledge when pressured by misleading context. This "calibration" process happens surprisingly fast, suggesting that the model’s internal representations already contain the necessary cues to distinguish between high-confidence parametric facts and low-quality external interference.
Table 6: $\beta$ value over training steps (ConFiQA-MC). Rapid convergence to a near-optimal regime that protects PK as fallback.
Step $\beta$
1 0.23
2 0.15
4 0.07
6 0.10
8 0.01
Appendix B Local vs. Global Advantage Weight Ablation
We ablate the local:global advantage combination ratio to confirm robustness to this hyperparameter. The "local" advantage ($A^{ck-local}$) focuses on optimizing the policy relative to the average of its own rollouts for a specific query-context pair, encouraging the model to find the best reasoning path within that specific evidence. In contrast, the "global" advantage ($A^{ck-global}$) penalizes the model based on its performance across the broader distribution of samples. Results on ConFiQA (Qwen2.5-7B) show that the 1:1 combination is near-optimal. This suggests that the model benefits equally from source-specific optimization and overall distributional stability. The fact that performance remains stable across the 1:1 and 1:2 ratios indicates that Knowledgeable-R1 is not overly sensitive to the exact weighting of these two signals, provided both are present to balance specialized reasoning with general robustness.
Table 7: Local:global advantage ratio ablation. Combining both is better than either alone; the exact ratio matters little.
Local:Global Setting Avg. score
1:0 local-only 60.95
0:1 global-only 64.69
1:1 local+global 65.40
1:2 local+global 65.20
Appendix C S1 Performance Gap and Component Weighting
A key discussion arises regarding the performance gap on S1 (correct context) compared to the standard GRPO w/ RAG baseline. To investigate this, we performed a sweep over the weights of our three objectives ($J_{\mathrm{PK}},J_{\mathrm{CK}},J_{\mathrm{RPK}}$). As shown in Table 8, the S1 performance is extremely sensitive to the weight of the contextual objective $J_{\mathrm{CK}}$. By increasing its relative weight to a 1:2:1 ratio, we can almost entirely close the gap with the baseline (78.04% vs. 77.36%) while still maintaining a massive lead in the adversarial S2 scenario (32.73% vs. 19.74%). This confirms that there is a tunable "Pareto front" between absolute mastery of reliable context and robustness to unreliable context. We chose the 1:1:1 ratio as our default because it offers a more conservative "safety first" profile, prioritizing the prevention of catastrophic failures in adversarial settings.
Table 8: Ablation of $J_{\mathrm{PK}}$$J_{\mathrm{CK}}$$J_{\mathrm{RPK}}$ weights on ConFiQA-MC (Qwen2.5-7B). Higher $J_{\mathrm{CK}}$ weight closes the S1 gap.
$J_{\mathrm{PK}}$$J_{\mathrm{CK}}$$J_{\mathrm{RPK}}$ S1 (correct ctx) S2 (adversarial)
1:2:1 78.04% 32.73%
1:1:1 (ours) 75.51% 37.34%
0:1:1 75.68% 35.53%
0:1:0 (GRPO w/ RAG) 77.36% 19.74%
Appendix D Query-only vs. Query+Context Behavior
[…]
Table 9: Accuracy under query-only vs. query+context inference (ConFiQA-MC, Qwen2.5-7B). Our method benefits from context when it is correct and resists it when wrong.
Method S1 (correct ctx) S2 (wrong ctx) Avg.
Query-only (base) 27.72% 25.92% 26.82%
RAG prompting 65.68% 13.47% 39.58%
GRPO-RAG (query-only) 32.48% 31.48% 31.98%
GRPO-RAG (context) 77.56% 26.94% 52.25%
Ours (query-only) 35.64% 36.70% 36.17%
Ours (context) 75.08% 43.94% 59.51%
Appendix E Sensitivity to Good/Bad Context Ratio
A common concern in RL for RAG is whether the model just learns the "noise ratio" of the training data. To test this, we trained Knowledgeable-R1 on data where bad contexts were extremely rare (only 1%). As Table 10 shows, Knowledgeable-R1 still manages to improve S2 performance significantly compared to GRPO w/ RAG. This is a crucial finding: it suggests that Knowledgeable-R1 is not just learning a naive prior that says "the context is often wrong," but is instead learning a genuine decision boundary that triggers parametric knowledge whenever the context is perceived as unreliable. Even a small number of adversarial signals during training are sufficient to ground the model’s reasoning process in its own parametric knowledge.
Table 10: Sensitivity to good:bad context ratio in training (ConFiQA-MR, Qwen2.5-7B). Even with 99% good context, Knowledgeable-R1 improves S2 resistance.
Good:Bad Ratio Method S1 (correct ctx) S2 (wrong ctx)
99%:1% RAG 65.68% 13.47%
GRPO w/ RAG 73.60% 14.31%
Knowledgeable-R1 72.11% 17.34%
75%:25% RAG 65.68% 13.47%
GRPO w/ RAG 74.92% 20.03%
Knowledgeable-R1 73.76% 29.29%
50%:50% (default) RAG 65.68% 13.47%
GRPO w/ RAG 77.56% 26.94%
Knowledgeable-R1 75.08% 43.94%
Appendix F Partial-Conflict Sensitivity
We construct test sets varying the number of misleading passages (1, 2, 4, 8) while keeping the query and gold answer fixed. Knowledgeable-R1 consistently outperforms GRPO w/ RAG across all levels of context adversariality.
Table 11: Accuracy vs. number of misleading passages. Knowledgeable-R1 maintains a substantial lead across all levels. (ConFiQA-MC, Qwen2.5-7B)
# misleading passages 1 2 4 8
GRPO w/ RAG 26.94% 26.60% 25.76% 25.42%
Knowledgeable-R1 (ours) 43.94% 43.10% 42.26% 40.07%
Appendix G Context-Dependence Diagnostics
To verify that robustness arises from the learned PK/CK arbitration rather than coincidence, we evaluate on subsets stratified by knowledge correctness and report two explicit diagnostic metrics on a test set where the context is designed to be wrong:
  • Factuality-score (w.r.t. context): This metric measures the linguistic and semantic alignment between the model’s answer and the provided context. A high score when context is wrong indicates that the model has been "led astray" by the evidence. Lower is better in this scenario.
  • KL-score: This measures the KL divergence between the output probability distributions generated with vs. without context. It acts as a structural measure of "contextual reliance." If the KL-score is low, the model is essentially ignoring the context and relying on its parametric prior. Lower is better when context is known to be misleading.
As seen in Table 12, Knowledgeable-R1 achieves significantly lower scores on both metrics, proving it has learned to actively decouple its reasoning from unreliable external inputs.
Table 12: Context-dependence diagnostics on wrong-context test set (ConFiQA-MC). Lower factuality-score and KL-score confirm Knowledgeable-R1 relies less on misleading context.
Method Factuality-score $\downarrow$ KL-score $\downarrow$
GRPO w/ RAG 0.40 213.73
Knowledgeable-R1 (ours) 0.16 168.83
To further rule out that robustness is purely inherited from the base model, we stratify by whether RAG prompting gives the correct answer (Table 13). On the subset where RAG prompting always fails, our method achieves 32.72%, far ahead of SFT (18.73%) and GRPO (19.92%), indicating the improvement stems from the learned PK/CK arbitration mechanism.
Table 13: Accuracy stratified by RAG-prompting outcome (ConFiQA-MC, wrong-context test set). Our method gains the most on the hardest subset where RAG prompting always fails.
Method RAG-answer correct $\uparrow$ RAG-answer wrong $\uparrow$
RAG prompting 100% 0%
SFT 90.27% 18.73%
GRPO w/ RAG 96.61% 19.92%
Knowledgeable-R1 (ours) 95.25% 32.72%
Appendix H Prompt designed in Knowledgeable-R1 and baselines